OpenStudy (anonymous):

Mrs. Chen has three dogs: one small, one medium, and one large. She buys dog bones at Jay's, where small bones cost 8 cents each, medium bones cost 13 cents and large bones cost 19 cents each. Mrs. Chen spent 85 cents for 6 dog bones. If she bought all three sizes, how many bones of each size did she buy? please show work

OpenStudy (anonymous):

# of small bones = X # of medium bones = Y # of large bones = Z X+Y+Z = 6 8X+13Y+19z = 85 @RagingSquirrel Please don't give out answers w/o work, especially when it's wrong (8+13+19 = 40, 40x2=80, not 85)

OpenStudy (anonymous):

yes, it is not 85

OpenStudy (ragingsquirrel):

My bad...